OpenStudy (anonymous):

(-4z³ + 9d) - (-12z³ + 6d) =

OpenStudy (anonymous):

Distribute the negative (subtraction) into ().

OpenStudy (anonymous):

-4z^3+9d+12z^3-6d =8z^3+3d

OpenStudy (anonymous):

I came up with -16z³ - 3d. This is wrong isn't it?

OpenStudy (anonymous):

(-)*(-)=(+) (-)*(+)=(-)

OpenStudy (anonymous):

-4z^3+9d+12z^3-6d simplify by combining like terms

OpenStudy (anonymous):

yes its wrong...see 12z^3 -4z^3 is 8z^3

OpenStudy (anonymous):

negative multiplied by a negative becomes a positive

OpenStudy (anonymous):

I'm so confused. Why are you multiplying?

OpenStudy (anonymous):

positive multiplied by a negative is negative.u have o remeber dis.then u can easily solve.

OpenStudy (anonymous):

u have to open the brackets...so u have to multiply

OpenStudy (anonymous):

A subtraction sign in front of ( ) is really multiplying everything in the ( ) by -1

OpenStudy (anonymous):

8z^3+3 d